1. A cement factory recently lowered its price from $60 to $40. As it did, the number of cement bags sold increased from 240000 to 280000.



I) what is the cement factory elasticity of demand? Interpret the result intuitively.



II) given that elasticity, did it make sense for the cement factory to lower its price?

1
Expert's answer
2022-05-10T13:22:25-0400

"\eta=\frac{\frac{60-40}{60+40}}{\frac{280000-240000}{280000+240000}}=2.6>1," elastic demand.